# Contrastive Sequential-Diffusion Learning: Non-linear and Multi-Scene Instructional Video Synthesis

### v2

@@ 5 Experimental Setting
In this section, we describe the experimental setup used to evaluate the performance of CoSeD in generating multi-scene video and image sequences for manual tasks. We provide details about the dataset used, the backbone models employed, and the baselines chosen for comparison. The aim is to demonstrate CoSeD’s ability to generalize across different models and generate coherent task-oriented outputs.
@@
We used a dataset [5] consisting of publicly available manual tasks in the recipes and DIY domains. Each manual task has a title, a description, a list of ingredients/resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated. Details about the dataset are in Annex A.
@@
Since CoSeD is independent of any specific video generation approach, we experimented with both Stable Video Diffusion [3] and Lumiere [2] models for multi-scene video generation. Stable Video Diffusion was selected for its public availability, while Lumiere was chosen for its capability to represent motion effectively.
@@
To evaluate CoSeD’s effectiveness in generating coherent image and video sequences for real-world manual tasks, we compared its performance against existing approaches: TALC [1] with ModelScope [28] and with Lumiere [2], SD 2.1 [24] with Stable Video Diffusion [3], stand-alone Lumiere [2], and for image sequences we tested Gill [14] and Seed-LLama [10]. During the evaluation, we prompted all models to generate a complete task.
@@ 6.1 Human evaluation
The human evaluation was conducted with a focus on two primary criteria. Annotators were asked to assess Semantic Consistency, which measures how well the video matches the instructional text, and Sequence Consistency, which involves rating the video on text alignment, visual consistency and overall quality of the video. See Annex C for details.
@@
The results in Table 1 show that CoSeD combined with Lumiere achieves the highest Sequence Consistency at 74.2% and leads in Semantic Consistency with 85.0%, highlighting its superior performance. This demonstrates that the combination of CoSeD and Lumiere is particularly effective for multi-scene generation tasks.
In contrast, methods like TALC + ModelScope and TALC + Lumiere show significantly lower Semantic Consistency scores (38.3% and 30.0%, respectively) and a Sequence Consistency of 50.8%. Altough SD + SVD and Lumiere alone perform better, they still do not match the performance of CoSeD + Lumiere, underscoring the advantages of our approach in achieving better coherence and text adherence in the generated videos.
@@
Table 1: Manual evaluation of multi-scene video generation models based on two key criteria: Semantic Consistency, evaluating text alignment with the instructional plan, and Sequence Consistency, assessing visual coherence, text alignment, and overall video quality.
@@
We also report the length of the generated videos in Table 1. CoSeD-based methods (CoSeD + SVD and CoSeD + Lumiere) achieves an average video length of around 15 and 21 seconds, respectively. This is substantially longer than TALC-based methods (TALC + ModelScope and TALC + Lumiere) which generate shorter videos, around 7 and 5 seconds on average.
A visual inspection of the generated videos (Figure 7) clearly indicate that CoSeD successfully depict all steps in the task, whereas TALC, despite being a multi-scene model, cannot achieve it. Even when TALC successfully depicts all steps, each scene typically lasts no more than 1.5 seconds, considering the average task length of 4.9 scenes (as detailed in Annex A). In contrast, our model consistently achieves at least 3 seconds per scene, effectively providing double the duration for each step.
Side-by-Side Evaluation
To assess models prioritizing coherence against our best model, we conduct a side-by-side evaluation. Annotators choose which videos better represent task steps, directly comparing each model’s coherence across scenes. See Annex C for details. This evaluation focuses on our top model, CoSeD + Lumiere, against coherence-focused models such as TALC + ModelScope, TALC + Lumiere, CoSeD + SVD, and the second-best model from Table 1, Lumiere.
According to the side-by-side evaluation results, Figure 8, CoSeD + Lumiere consistently outperforms all other models, with annotators repeatedly selecting it over competitors. For instance, CoSeD + Lumiere achieves an 87% selection rate compared to just 13% for TALC + ModelScope, demonstrating its superior ability to maintain coherence. It is also chosen 68% of the time over TALC + Lumiere, which has a selection rate of 32%, reflecting its better task consistency.
@@
To evaluate the absolute quality of the generated video sequences, human annotators rated each sequence on a scale from 1 to 5, comparing them to ground-truth sequences. Deviations like hallucinated visual artifacts or inconsistent actions affect perceived quality. As shown in Table 2, our method achieves over 64% of the ground-truth score, with ground-truth sequences scoring just 0.5 points below the maximum.
@@ 6.2 Automatic evaluation
We employ CLIP [22] to evaluate the sequence similarity of each task ($V\mapsto V$) and to assess the adherence of the generated image to the given textual descriptions ($T\mapsto V$). This provides a comprehensive evaluation of the generated images alignment with the intended textual descriptions and their visual coherence across the entire sequence.
@@
CoSeD Performance
Our model achieves a sequence similarity score ($V\mapsto V$) of $84.8$ and a description adherence score ($T\mapsto V$) of approximately $27.1$ (see Table 3), outperforming the image and video baselines in textual adherence. These results highlight CoSeD’s ability to generate sequences that are both visually coherent and semantically aligned with the text, demonstrating its strong potential for practical applications.
Comparison with Baseline Models
In a comparison to other sequence-generating models (see Table 3), CoSeD with both Video Stable Diffusion and Lumiere consistently achieves the highest description adherence ($T\mapsto V$) without discarding the sequence similarity ($V\mapsto V$).
When comparing CoSeD with the best model in sequence similarity, our model lags by only 5.9% while gains 11.8% in description adherence. While TALC + Lumiere excels in maintaining high sequence similarity, CoSeD demonstrates superior description adherence without compromising sequence similarity. This strong description adherence score highlights our model’s effectiveness in aligning generated content with text descriptions, which is crucial for tasks such as accurately converting textual descriptions into videos. Compared to vanilla video-only models, our approach surpasses both metrics, leading to improved results.
@@ 6.3 Qualitative Analysis
Figure 7 and 7 provides a closer look at how different methods influence the quality of generated videos and image sequences for a certain recipe. Our model excels in maintaining a consistent background, keeping the same pan, and ensuring the ingredients evolve seamlessly from raw to final recipe. This accurate depiction of sequence provides a visually stable and easy-to-follow experience for viewers.
@@
Figure 10: The average number of times that CoSeD selected a given step to generate the next scene.
@@
For the out-of-scope DIY tasks shown in Figures 7 and 7, our model effectively handles broader actions, such as showing a room without furniture or the process of cleaning, with good text adherence. However, it struggles with depicting complex tools like vacuum cleaners. Other methods often generate very similar images, failing to accurately represent the task or omitting steps entirely, as demonstrated by TALC.
@@ 6.4 Ablation Studies
This section analysis CoSeD ’s contrastive selection of latents and steps, and its role in ensuring the alignment with (possibly non-linear) sequences of instructions.
@@
Denoising latents.
Understanding how latent variables impact sequence coherence and adherence is key to refining our model. Figure 11 shows a correlation between latent settings and the model’s ability to produce coherent, textually aligned sequences.
The analysis in Figure 11 identifies four key latent configuration areas. The ideal zone balances high sequence similarity (CLIP Image Score) and description adherence (CLIP Text Score). Latent 5 shows moderate text adherence but lacks visual coherence. Latents 20 and 40 produce coherent images but deviate from task steps, resulting in poor text adherence. Latent 10 performs poorly overall. Figure 12 provides examples from these four key latent areas: Latent 5 generates images that adhere to the prompt but lacks coherence, Latent 10 starts with good coherence and text alignment but loses coherence over time, and Latent 20 produces overly similar images with low text adherence.
Ultimately, CoSeD (top-right mark in Figure 11) achieves superior performance by leveraging early denoising latents, using contrastive selection to enhance results compared to using individual latents.
@@ 7 Conclusion
Generating multi-scene instructional videos for complex tasks, as DIY projects and recipes, all while maintaining sequence coherence and accurate scene representation, is not a trivial feat.
Our method addresses these challenges with key contributions. First, we employ a decoder to generate visual prompts from the sequence of instructions to ground the generation in a common context, thus, ensuring better scene accuracy. Second, by conditioning the diffusion process on images from previous scenes, the method maintains coherence across scenes. Third, and more importantly, the CoSeD’s contrastive selection of the most consistent image, enables the assessement of all previous steps. The result is the selection of the image that is most related with the overall sequence rather than just the preceding step. Also, the contrastive nature of CoSeD, allows generating non-linear sequences of video scenes – a unique property of CoSeD.
Additionally, the model’s compact design helps with achieving an efficient training and easy domain-specific fine-tuning, while its flexibility supports ensembles of diffusion models for optimal performance.
Evaluations confirm that our method effectively maintains scene coherence and accurately represents textual descriptions, as demonstrated by both manual and automatic evaluations. Importantly, side-by-side human annotation shows that annotators prefer our model over 65% of the time, highlighting the effectiveness of our sequence-grounded approach. This preference underscores the value of our method in producing high-quality, coherent instructional videos.
@@ Appendix A Dataset
Each task in the dataset includes a title, a description, a list of ingredients/resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated. To facilitate the illustration of task steps, we focused on tasks that are mostly illustrated, allowing us to use these images as ground truth for training and evaluating our methods.
The dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, amounting to a total of 6,860 individual steps. Most tasks include an image for each step, and some feature a complete recipe video that is segmented into multiple clips, with each clip lasting between 10 and 30 seconds per step.
Considering that the number of illustrations can impact accuracy, we limited the training tasks to those with no more than 10 steps.
@@ Appendix B Model Training
We opted for the CLIP model with a patch size of 32 to serve as the encoder for both image and text data due to its reputation in effectively capturing visual and textual information. In training our own architecture, we conducted experiments with various hyperparameters, including different learning rates, learning rate schedulers, dropout rates, layer freezing, and batch sizes, in order to identify the most suitable settings for our specific problem.
For the loss function, we employed cross-entropy, comparing the softmax output with hot-encoding of the steps that belong to each task. It’s important to note that while this loss function indicates step-task associations, it may not always accurately reflect the model’s overall performance on the task at hand. During the tunying of hyperparameters, we found that freezing layers, weight decay, dropout, and learning rate schedulers had minimal impact on model performance.
@@
The best model, which has about 600,000 parameters, was refined using specific training parameters listed in Table 4. Training was completed in under two minutes, utilizing an A100-40GB GPU and spanning across ten epochs. Employing the Cross-Entropy loss function, the training process operated with a batch size of 500 and a learning rate set at 0.01, using the Adam optimizer.
@@
Prompt Rewriter Training

### v3

@@ 5 Experimental Setting
In this section, we describe the experimental setup used to evaluate the performance of CoSeD in generating multi-scene video and image sequences for manual tasks. We provide details on the dataset used, the backbone models, and the baselines chosen for comparison. The aim is to demonstrate CoSeD’s ability to generalize across different models and generate coherent task-oriented outputs.
@@
We used a dataset [5] consisting of publicly available manual tasks in recipes and DIY domains. Each manual task has a title, a description, a list of ingredients, resources, and tools, and a sequence of step-by-step instructions, which may or may not be illustrated. Details about the dataset can be found in the appendix file.
@@
Since CoSeD is independent of the video generation method, we experimented with both Stable Video Diffusion [3] and Lumiere [2] models for multi-scene video generation. Stable Video Diffusion was selected for its public availability, while Lumiere was chosen for its enhanced capability to represent complex motion effectively.
@@
To evaluate the effectiveness of CoSeD in generating coherent image and video sequences for real-world manual tasks, we compared its performance against existing approaches: TALC [1] with ModelScope [28] and with Lumiere [2], SD 2.1 [24] with Stable Video Diffusion [3], stand-alone Lumiere [2], and for image sequences we tested Gill [14] and Seed-LLama [10]. During the evaluation, we prompted all models to generate a complete task.
@@ 6.1 Human evaluation
The human evaluation was conducted with a focus on two primary criteria. Annotators were asked to assess Semantic Consistency, which measures how well the video matches the instructional text, and Sequence Consistency, which involves rating the video on text alignment, visual consistency and video quality. See the appendix file for details.
@@
The results in Table 1 show that CoSeD combined with Lumiere achieves the highest Sequence Consistency at 74.2% and leads in Semantic Consistency with 85.0%, highlighting its effectiveness. This shows that the combination of CoSeD and Lumiere is particularly effective for multi-scene generation tasks.
In contrast, methods such as TALC + ModelScope and TALC + Lumiere show significantly lower Semantic Consistency scores (38.3% and 30.0%, respectively) and a Sequence Consistency of 50.8%. Although SD + SVD and Lumiere alone perform better, they still do not match the performance of CoSeD + Lumiere, underscoring the advantages of our approach in achieving better coherence and text adherence in the generated videos.
@@
Table 1: Manual evaluation of multi-scene video generation models based on two key criteria: Semantic Consistency, which measures the alignment of generated content with the described task steps, and Sequence Consistency, which assesses the visual coherence, text alignment, and overall quality of the video.
@@
We also report the length of the videos generated in Table 1. CoSeD-based methods (CoSeD + SVD and CoSeD + Lumiere) achieve an average video length of around 15 and 21 seconds, respectively. This is substantially longer than TALC-based methods (TALC + ModelScope and TALC + Lumiere) which generate shorter videos, around 7 and 5 seconds on average.
A visual inspection of the generated videos (Figure 7) clearly indicates that CoSeD successfully depicts all steps in the task, whereas TALC, despite being a multi-scene model, cannot achieve it. Even when TALC successfully depicts all steps, each scene typically lasts no more than 1.5 seconds, considering the average task length of 4.9 scenes (as detailed in the appendix file). In contrast, our model consistently achieves at least 3 seconds per scene, effectively providing double the duration for each step.
Side-by-Side Evaluation.
To assess models prioritizing coherence against our best model, we conduct a side-by-side evaluation. Annotators choose which videos better represent task steps, directly comparing each model’s coherence across scenes. See the appendix file for details. This evaluation focuses on our best model, CoSeD + Lumiere, against coherence-focused models such as TALC + ModelScope, TALC + Lumiere, CoSeD + SVD, and the second-best model from Table 1, Lumiere.
According to the side-by-side evaluation results, Figure 8, CoSeD + Lumiere consistently outperforms all other models, with annotators repeatedly selecting it over competitors. For example, CoSeD + Lumiere achieves a selection rate 87% compared to just 13% for TALC + ModelScope, demonstrating its superior ability to maintain coherence. It is also chosen 68% of the time over TALC + Lumiere, which has a selection rate of 32%, reflecting its better task consistency.
@@
To evaluate the absolute quality of the generated video sequences, human annotators rated each sequence on a scale of 1 to 5, comparing them to ground-truth sequences. Deviations like hallucinated visual artifacts or inconsistent actions affect perceived quality. As shown in Table 2, our method achieves more than 64% of the ground truth score, with ground truth sequences scoring just 0.5 points below the maximum.
@@ 6.2 Automatic evaluation
We employ CLIP [22] to evaluate the sequence similarity of each task ($V\mapsto V$) and to assess the adherence of the generated image to the given textual descriptions ($T\mapsto V$). This provides a comprehensive evaluation of the alignment of the generated images with the intended textual descriptions and their visual coherence throughout the sequence.
@@
CoSeD Performance.
CoSeD achieves a sequence similarity score ($V\mapsto V$) of $84.8$ and a description adherence score ($T\mapsto V$) of approximately $27.1$ (see Table 3), outperforming the image and video baselines in textual adherence. These results highlight CoSeD’s ability to generate sequences that are both visually coherent and semantically aligned with the text, demonstrating its significant potential for practical applications.
Comparison with Baseline Models.
In a comparison to other sequence-generating models (see Table 3), CoSeD with both Video Stable Diffusion and Lumiere consistently achieves the highest description adherence ($T\mapsto V$) without discarding sequence similarity ($V\mapsto V$).
When comparing CoSeD with the best model in sequence similarity, our model lags only $5.9\%$ while it gains $11.8\%$ in description adherence. Although TALC + Lumiere excels in maintaining high sequence similarity, CoSeD demonstrates superior adherence to descriptions without compromising sequence similarity. This strong description adherence score highlights the effectiveness of our model in aligning generated content with text descriptions, which is crucial for tasks such as accurately converting textual descriptions into videos. Compared to vanilla video-only models, our approach surpasses both metrics, leading to improved results.
@@ 6.3 Qualitative Analysis
Figure 7 and 7 provide a closer look at how different methods influence the quality of generated videos and image sequences for a certain recipe. Our model excels in maintaining a consistent background, keeping the same pan, and ensuring that the ingredients evolve seamlessly from raw to final recipe. This clear depiction of the sequence provides viewers with a visually stable and easy-to-follow video.
@@
Figure 10: The average number of times that CoSeD selected each task step for generating the next scene.
@@
For the out-of-scope DIY tasks shown in Figures 7 and 7, our model effectively handles broader actions, such as showing a room without furniture or a cleaning process, with good text adherence. However, it struggles with depicting complex tools such as vacuum cleaners. Other methods often generate very similar images, failing to accurately represent the task or omitting steps entirely, as demonstrated by TALC.
@@ 6.4 Ablation Studies
This section analysis CoSeD contrastive selection of latents and steps along with its role in ensuring alignment with (possibly non-linear) sequences of instructions.
Denoising Latents.
Understanding how latent variables affect sequence coherence and adherence is key to refining our model. Figure 11 shows a correlation between latent settings and the model’s ability to produce coherent, textually aligned sequences. The analysis in Figure 11 identifies four key latent configuration areas. The ideal zone balances high sequence similarity (CLIP Image Score) and description adherence (CLIP Text Score). Latent 5 shows moderate text adherence, but lacks visual coherence. Latents 20 and 40 produce coherent images but deviate from task steps, leading to weak text adherence. Latent 10 performs poorly overall.
@@
Figure 12 provides examples from these key latent areas: Latent 5 generates images that adhere to the prompt but lack coherence, Latent 10 starts with good coherence and text alignment but loses coherence over time, and Latent 20 produces overly similar images with low text adherence.
Ultimately, CoSeD (top right mark in Figure 11) achieves superior performance by leveraging early denoising latents, using contrastive selection to enhance results compared to using individual latents.
@@ 7 Conclusion
Generating multi-scene instructional videos for complex tasks, such as DIY projects and recipes, all while maintaining sequence coherence and accurate scene representation, is not a trivial feat. The proposed method addresses these challenges with key contributions. First, we employ a decoder model to generate visual prompts from the sequence of instructions to ground the generation in a common context, thus ensuring better scene accuracy. Second, by conditioning the diffusion process on images from previous scenes, the method maintains coherence across scenes. Third, and more importantly, the CoSeD’s contrastive selection of the most consistent image enables the assessment of all previous steps. The result is the selection of the image that is most related to the overall sequence rather than just the preceding step. Additionally, the contrastive nature of CoSeD enables the generation of non-linear sequences of video scenes, an exclusive feature of CoSeD.
In addition, the compact design of the model helps to achieve efficient training and easy domain-specific fine-tuning, while its flexibility supports ensembles of diffusion models for optimal performance.
Evaluations confirm that our method effectively maintains scene coherence and accurately represents textual descriptions, as demonstrated by both manual and automatic evaluations. Importantly, side-by-side human annotation shows that annotators prefer our model over $65\%$ of the time, highlighting the effectiveness of our sequence-grounded approach. This preference underscores the value of our method in producing coherent and high-quality instructional videos.
Acknowledgements
We thank anonymous reviewers for their valuable comments and suggestions. This work was partially supported by a Google Research Gift and by the FCT project NOVA LINCS Ref. (UIDB/04516/2020).
@@ Appendix A Dataset
Each task in the dataset includes a title, a description, a list of ingredients/resources and tools, and a sequence of step-by-step instructions, which may or may not be illustrated. To facilitate the illustration of task steps, we focused on tasks that are mostly illustrated, allowing us to use these images as the ground truth for training and evaluating our methods.
The dataset comprises approximately 1,400 tasks, with an average of 4.9 steps per task, which is a total of 6,860 individual steps. Most tasks include an image for each step, and some feature a complete recipe video that is segmented into multiple clips, with each clip lasting between 10 and 30 seconds per step.
Considering that the number of illustrations can affect the accuracy, we limited the training tasks to those with no more than 10 steps.
@@ Appendix B Model Training
We opted for the CLIP model with a patch size of 32 to serve as the encoder for both image and text data due to its reputation in effectively capturing visual and textual information. In training our own architecture, we conducted experiments with various hyperparameters, including different learning rates, learning rate schedulers, dropout rates, layer freezing, and batch sizes, to identify the most suitable settings for our specific problem.
For the loss function, we employed cross-entropy, comparing the softmax output with hot-encoding of the steps that belong to each task. It is important to note that while this loss function indicates step-task associations, it may not always accurately reflect the model’s overall performance on the task at hand. During the tuning of hyperparameters, we found that freezing layers, weight decay, dropout, and learning rate schedulers had minimal impact on model performance.
@@
The best model, which has about 600,000 parameters, was refined using specific training parameters listed in Table 4. Training was completed in under two minutes, using an A100-40GB GPU and spanning ten epochs. Employing the Cross-Entropy loss function, the training process operated with a batch size of 500 and a learning rate set at 0.01, using the Adam optimizer.
@@
Prompt Rewriter Training.